# Robust Bernoulli mixture models for credit portfolio risk

### v1

@@ Acknowledgements
The property that $C_{D_{n},Z}=\hat{C}_{X_{n},Y}$ (which is SI) follows from Remark 3.4(b). ∎
Proof of Proposition 4.1:.
By assumption, the random variables $X_{1},\ldots,X_{N}$ are conditionally independent given $Y$ and the random variables $X_{1}^{\prime},\ldots,X_{N}^{\prime}$ are conditionally independent given $Y^{\prime}\,.$ Hence, we obtain from (Ansari and Rüschendorf, 2024, Corollary 4) that $(X_{1},\ldots,X_{N})\leq_{sm}(X_{1}^{\prime},\ldots,X_{N}^{\prime})\,,$ using that $X_{n}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{n}^{\prime}$ for all $n$ and $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and using that the copulas $C_{X_{n},Y}$ and $C_{X_{n}^{\prime},Y^{\prime}}$ are SI and satisfy $C_{X_{n},Y}(u,v)\leq C_{X_{n}^{\prime},Y^{\prime}}(u,v)$ for all $(u,v)\in[0,1]^{2}\,,$ where $\leq_{sm}$ denotes the supermodular order which compares expectations of supermodular functions of random vectors. Since the supermodular order is closed under decreasing transformations ((Shaked and Shantikumar, 2007, Theorem 9.A.9(a))), it follows that $(D_{1},\ldots,D_{N})\leq_{sm}(D_{1}^{\prime},\ldots,D_{N}^{\prime})\,.$ Applying (Shaked and Shantikumar, 2007, Corollary 9.A.10), we obtain $(l_{1}D_{1},\ldots,l_{N}D_{N})\leq_{sm}(l_{1}D_{1}^{\prime},\ldots,l_{N}D_{N}^{\prime})$ using non-negativity as well as independence of $\{l_{n}\}_{n}$ from $\{D_{n}\}_{n}$ and $\{D_{n}^{\prime}\}_{n}\,.$ Since for any convex function $\varphi\colon\mathbb{R}\to\mathbb{R}\,,$ the function $(x_{1},\ldots,x_{N})\mapsto\varphi(x_{1}+\cdots+x_{N})$ is a supermodular function, see (Müller and Stoyan, 2002, Theorem 8.3.3 and its proof), the convex comparison in (4.1) follows. ∎
@@
We construct copula-based threshold models which coincide with the siBMMs and are built by SI copulas that allow a pointwise comparison. Then, Proposition 4.1 implies the statement.
To this end, consider $\{X_{n}\}_{n}$ and $Y$ defined in (5.25). We obtain from Lemma 3.5 that $X_{n},Y\sim U(0,1)$ and $C_{X_{n},Y}$ is SI for all $n\,.$ Further, $X_{1},\ldots,X_{N}$ are conditionally independent given $Y$ and we have $D_{n}=\mathds{1}_{\{X_{n}\leq c_{n}\}}$ for all $n\,.$ Similarly, consider random variables $(X_{n}^{\prime})_{n}$ and $Y^{\prime}$ associated with the default model $(D_{1}^{\prime},\ldots,D_{N}^{\prime},Z^{\prime})\,.$ Then, we obtain for all $(u,v)\in[0,1]^{2}$ and for all $n$ that
$C_{X_{n},Y}(u,v)$ $=\int_{0}^{v}P(X_{n}\leq u\mid Y=y)\mathrm{\,d}y$
$=1-\int_{v}^{1}P(X_{n}\leq u\mid Y=y)\mathrm{\,d}y$
=$=\begin{cases}1-\int_{v}^{1}\frac{u}{c_{n}}P(D_{n}=1\mid Z=F_{Z}^{-1}(1-y))\mathrm{\,d}y\,,&\text{if }u\leq c_{n}\,,\\ 1-\int_{v}^{1}\left[\frac{u-c_{n}}{1-c_{n}}+\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)P(D_{n}=1\mid Z=F_{Z}^{-1}(1-y))\right]\mathrm{\,d}y\,,&\text{if }u>c_{n}\,,\end{cases}$
=$=\begin{cases}1-\frac{u}{c_{n}}\int_{0}^{1-v}P(D_{n}=1\mid Z=F_{Z}^{-1}(y^{\prime}))\mathrm{\,d}y^{\prime}\,,&\text{if }u\leq c_{n}\,,\\ 1-\frac{(1-v)(u-c_{n})}{1-c_{n}}-\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)\int_{0}^{1-v}P(D_{n}=1\mid Z=F_{Z}^{-1}(y^{\prime}))\mathrm{\,d}y^{\prime}\,,&\text{if }u>c_{n}\,,\end{cases}$
=$=\begin{cases}1-\frac{u}{c_{n}}G_{n,Z}(1-v)\,,&\text{if }u\leq c_{n}\,,\\ 1-\frac{(1-v)(u-c_{n})}{1-c_{n}}-\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)G_{D_{n},Z}(1-v)\,,&\text{if }u>c_{n}\,,\end{cases}$
$\leq\begin{cases}1-\frac{u}{c_{n}}G_{n,Z^{\prime}}(1-v)\,,&\text{if }u\leq c_{n}\,,\\ 1-\frac{(1-v)(u-c_{n})}{1-c_{n}}-\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)G_{D_{n},Z^{\prime}}(1-v)\,,&\text{if }u>c_{n}\,,\end{cases}$
@@
where the third equality follows from (5.30). For the fourth equality, we transform $y^{\prime}:=1-y\,.$ The fifth equality holds true by definition of $G_{n,Z}$ in (2.2), and the inequality is due to the assumption that $G_{n,Z}(z)\geq G_{n,Z^{*}}(z)$ for all $z\,,$ where we use that $\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\geq 0\,.$
@@
$F_{D_{n}|Z=F_{Z}^{-1}(\cdot)}(y)=_{S}F_{D_{n}^{\prime}|Z^{\prime}=F_{Z^{\prime}}^{-1}(\cdot)}(y)\quad\text{for all }y\in\mathbb{R}\,,$ (5.32)
where $=_{S}$ denotes equality in the Schur order for functions, i.e., the increasing rearrangements of the left and right sides of (5.32) coincide. Due to the assumptions, we know that $D_{n}^{\prime}$ is stochastically increasing in $Z^{\prime}\,,$ and $D_{n}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}D_{n}^{\prime}$ for all $n$ as well as $Z\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z^{\prime}\,.$ Hence, we obtain from (Ansari and Ritter, 2024, Proposition 5.4) that $(D_{1},\ldots,D_{N})\leq_{sm}(D_{1}^{\prime},\ldots,D_{N}^{\prime})\,.$ Then, proceeding in the same way as in the end of the proof of Proposition 4.1, we obtain (4.2). ∎
@@
Since $G_{D_{n}^{x},Z^{x}}\in\mathcal{F}_{icx}^{\pi_{n}}$ for all $x\in I$ and since, for any fixed $\alpha\,,$ the class $\mathcal{F}_{icx}^{\alpha}$ is closed under pointwise minima and maxima, it follows that $\underline{G}_{n},\overline{G}_{n}\in\mathcal{F}_{icx}^{\pi_{n}}\,.$ Then, the assertions follow from Theorem 2.1 using the one-to-one correspondence between default integral functions and the class $\mathcal{F}_{icx}^{\pi_{n}}\,.$

### v2

@@ Appendix B Proofs
The property that $C_{D_{n},Z}=\hat{C}_{X_{n},Y}$ (which is SI) follows from Remark A.3(b). ∎
Proof of Proposition A.6:.
By assumption, the random variables $X_{1},\ldots,X_{N}$ are conditionally independent given $Y$ and the random variables $X_{1}^{\prime},\ldots,X_{N}^{\prime}$ are conditionally independent given $Y^{\prime}.$ Hence, we obtain from (Ansari and Rüschendorf, 2024, Corollary 4) that $(X_{1},\ldots,X_{N})\leq_{sm}(X_{1}^{\prime},\ldots,X_{N}^{\prime}),$ using that $X_{n}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X_{n}^{\prime}$ for all $n$ and $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and using that the copulas $C_{X_{n},Y}$ and $C_{X_{n}^{\prime},Y^{\prime}}$ are SI and satisfy $C_{X_{n},Y}(u,v)\leq C_{X_{n}^{\prime},Y^{\prime}}(u,v)$ for all $(u,v)\in[0,1]^{2},$ where $\leq_{sm}$ denotes the supermodular order which compares expectations of supermodular functions of random vectors. Since the supermodular order is closed under decreasing transformations ((Shaked and Shantikumar, 2007, Theorem 9.A.9(a))), it follows that $(D_{1},\ldots,D_{N})\leq_{sm}(D_{1}^{\prime},\ldots,D_{N}^{\prime}).$ Applying (Shaked and Shantikumar, 2007, Corollary 9.A.10), we obtain $(l_{1}D_{1},\ldots,l_{N}D_{N})\leq_{sm}(l_{1}D_{1}^{\prime},\ldots,l_{N}D_{N}^{\prime})$ using non-negativity as well as independence of $\{l_{n}\}_{n}$ from $\{D_{n}\}_{n}$ and $\{D_{n}^{\prime}\}_{n}.$ Since for any convex function $\varphi\colon\mathbb{R}\to\mathbb{R}$, the function $(x_{1},\ldots,x_{N})\mapsto\varphi(x_{1}+\cdots+x_{N})$ is a supermodular function, see (Müller and Stoyan, 2002, Theorem 8.3.3 and its proof), the convex comparison in (A.11) follows. ∎
@@
We construct copula-based threshold models which coincide with the siBMMs and are built by SI copulas that allow a pointwise comparison. Then, Proposition A.6 implies the statement.
To this end, consider $\{X_{n}\}_{n}$ and $Y$ defined in (B.2). We obtain from Lemma A.4 that $X_{n},Y\sim U(0,1)$ and $C_{X_{n},Y}$ is SI for all $n.$ Further, $X_{1},\ldots,X_{N}$ are conditionally independent given $Y$ and we have $D_{n}=\mathds{1}_{\{X_{n}\leq c_{n}\}}$ for all $n.$ Similarly, consider random variables $(X_{n}^{\prime})_{n}$ and $Y^{\prime}$ associated with the default model $(D_{1}^{\prime},\ldots,D_{N}^{\prime},Z^{\prime}).$ Then, we obtain for all $(u,v)\in[0,1]^{2}$ and for all $n$ that
$C_{X_{n},Y}(u,v)$ $=\int_{0}^{v}\mathbb{P}(X_{n}\leq u\mid Y=y)\mathrm{\,d}y=u-\int_{v}^{1}\mathbb{P}(X_{n}\leq u\mid Y=y)\mathrm{\,d}y$
$=\begin{cases}u-\int_{v}^{1}\frac{u}{c_{n}}\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(1-y))\mathrm{\,d}y,&\text{if }u\leq c_{n},\\ u-\int_{v}^{1}\left[\frac{u-c_{n}}{1-c_{n}}+\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(1-y))\right]\mathrm{\,d}y,&\text{if }u>c_{n},\end{cases}$
$=\begin{cases}u-\frac{u}{c_{n}}\int_{0}^{1-v}\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(y^{\prime}))\mathrm{\,d}y^{\prime},&\text{if }u\leq c_{n},\\ u-\frac{(1-v)(u-c_{n})}{1-c_{n}}-\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)\int_{0}^{1-v}\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(y^{\prime}))\mathrm{\,d}y^{\prime},&\text{if }u>c_{n},\end{cases}$
$=\begin{cases}u-\frac{u}{c_{n}}G_{n,Z}(1-v),&\text{if }u\leq c_{n},\\ u-\frac{(1-v)(u-c_{n})}{1-c_{n}}-\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)G_{D_{n},Z}(1-v),&\text{if }u>c_{n},\end{cases}$
$\leq\begin{cases}u-\frac{u}{c_{n}}G_{n,Z^{\prime}}(1-v),&\text{if }u\leq c_{n},\\ u-\frac{(1-v)(u-c_{n})}{1-c_{n}}-\left(\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\right)G_{D_{n},Z^{\prime}}(1-v),&\text{if }u>c_{n},\end{cases}$
@@
where the third equality follows from (B.7). For the fourth equality, we transform $y^{\prime}:=1-y.$ The fifth equality holds true by definition of $G_{n,Z}$ in (2.2), and the inequality is due to the assumption that $G_{n,Z}(z)\geq G_{n,Z^{*}}(z)$ for all $z,$ where we use that $\frac{u}{c_{n}}-\frac{u-c_{n}}{1-c_{n}}\geq 0.$
@@
$F_{D_{n}|Z=F_{Z}^{-1}(\cdot)}(y)=_{S}F_{D_{n}^{\prime}|Z^{\prime}=F_{Z^{\prime}}^{-1}(\cdot)}(y)\quad\text{for all }y\in\mathbb{R},$ (B.8)
where $_{S}$ denotes equality in the Schur order for functions, i.e., the increasing rearrangements of the left and right sides of (B.8) coincide. Due to the assumptions, we know that $D_{n}^{\prime}$ is stochastically increasing in $Z^{\prime},$ and $D_{n}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}D_{n}^{\prime}$ for all $n$ as well as $Z\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Z^{\prime}.$ Hence, we obtain from (Ansari and Ritter, 2025, Proposition 4.2) that $(D_{1},\ldots,D_{N})\leq_{sm}(D_{1}^{\prime},\ldots,D_{N}^{\prime}).$ Then, proceeding in the same way as in the end of the proof of Proposition A.6, we obtain (2.4). ∎
@@
Using $G_{D_{n}^{x},Z^{x}}\in\mathcal{F}_{icx}^{\pi_{n}}$ for all $x\in I$, it follows that $\underline{G}_{n},\overline{G}_{n}\in\mathcal{F}_{icx}^{\pi_{n}}$ with $\underline{G}_{n}\geq G_{D_{n}^{x},Z^{x}}\geq\overline{G}_{n}$ pointwise for all $x.$ Then, the assertions follow from Theorem 2.1 using the one-to-one correspondence between default integral functions and the class $\mathcal{F}_{icx}^{\pi_{n}}.$